chocolateluva77:

the sum of two numbers is 12. the sum of triple the larger number and four times the smaller number is 40. setup and solve a system of equations to find the nunbers. pls explain how u got the steps to the answer. ty.

chocolateluva77:

*numbers

GabeYoungEpic2008:

umm all i can think of is 9 and 3 bc 9x3=27 3x4=12 27+12=39 i hope this helps

GabeYoungEpic2008:

9 and 3.25 is the right answer 9x3=27 3.25x4=13 27+13=40

GabeYoungEpic2008:

please mark it a medal

chocolateluva77:

where did u get 3.25 from?

GabeYoungEpic2008:

it says 2 numbers so 9 and 3.25

GabeYoungEpic2008:

i was just random thinking

chocolateluva77:

but 9 and 3.25 doesn't equal 12

GabeYoungEpic2008:

oh yea sorry ill rethink

chocolateluva77:

okay

GabeYoungEpic2008:

its 12 and 1 12x3= 36 1x4=4 36+4=40

GabeYoungEpic2008:

thats it i earn a medal

chocolateluva77:

12 and 1 doesn't equal 12

Vocaloid:

The problem says to set up a system of equations. Let’s let our two numbers be x and y. Let’s let x be the smaller one and y be the larger one. the sum of two numbers is 12 —> x + y = 12 the sum of triple the larger number and four times the smaller number is 40 —> triple the larger one = 3y, and four times the smaller one = 4x. Their sum is 40. The second equation is therefore 4x + 3y = 40 So our system now is x + y = 12 4x + 3y = 40 Solve using substitution, elimination, or graphing.

chocolateluva77:

ty